x^2+4x+4=7
How do I solve this?

Answer by fractalier(6550)

A couple of ways...let me show you how to do it without using the quadratic formula...notice the left side is a perfect square...let's take advantage of that...we get
x^2 + 4x + 4 = 7
(x + 2)^2 = 7
x + 2 = ± sqrt(7)
x = -2 ± sqrt(7)
and we're done...